### v1

# A Two-Step Procedure to Detect Cosmological Gravitational Wave Backgrounds with Next-Generation Terrestrial Gravitational-Wave Detectors

Abstract
Cosmological gravitational-wave backgrounds are an exciting science target for next-generation ground-based detectors, as they encode invaluable information about the primordial Universe. However, any such background is expected to be obscured by the astrophysical foreground from compact-binary coalescences. We propose a novel framework to detect a cosmological gravitational-wave background in the presence of binary black holes and binary neutron star signals with next-generation ground-based detectors, including Cosmic Explorer and the Einstein Telescope. Our procedure involves first removing all the individually resolved binary black hole signals by notching them out in the time-frequency domain. Then, we perform joint Bayesian inference on the individually resolved binary neutron star signals, the unresolved binary neutron star foreground, and the cosmological background. For a flat cosmological background, we find that we can claim detection at $5\,\sigma$ level when $\Omega_{\mathrm{ref}}\geqslant 2.7\times 10^{-12}/\sqrt{T_{\mathrm{obs}}/\mathrm{yr}}$, where $T_{\mathrm{obs}}$ is the observation time (in years), which is within a factor of $\lesssim 2$ from the sensitivity reached in absence of these astrophysical foregrounds.
preprint: FERMILAB-PUB-24-0871-T
GW
gravitational-wave
PSD
power spectral density
GR
general relativity
CBC
compact binary coalescence
BH
black hole
BBH
binary black hole
BNS
binary neutron star
NSBH
neutron star-black hole
SFR
star formation rate
LVK
LIGO-Virgo-KAGRA
ET
Einstein Telescope
CE
Cosmic Explorer
PE
parameter estimation
SGWB
stochastic gravitational wave background
CGWB
cosmological gravitational wave background
XG
next generation
ML
Machine Learning
DL
Deep Learning
NN
Neural Network
SNR
signal-to-noise ratio
PDF
probability density function
CDF
cumulative density function
Introduction. The next generation (XG) of ground-based gravitational-wave (GW) detectors, such as Cosmic Explorer (CEReitze et al. (2019); Evans et al. (2021); Srivastava et al. (2022) and the Einstein Telescope (ETPunturo et al. (2010); Branchesi et al. (2023a), will see an unprecedented increase in sensitivity both to individual compact binary coalescence (CBC) events and to stochastic signals. This will allow us to individually resolve nearly all of the binary black hole (BBH) mergers and a significant fraction of binary neutron star (BNS) and neutron star-black hole (NSBH) events in the entire Universe Gupta et al. (2023); Branchesi et al. (2023b). Furthermore, we will probe the stochastic gravitational wave backgrounds across several orders of magnitude in amplitude Regimbau et al. (2017); Renzini et al. (2022). In particular, the detection of a cosmological gravitational wave background (CGWB) would open up a unique window to observe the earliest moments of the Universe and probe physics at energies close to the Planck scale Grishchuk (1974); Starobinsky (1979); Grishchuk (1993); Barnaby et al. (2012); Damour and Vilenkin (2005); Siemens et al. (2007); Olmez et al. (2010); Regimbau et al. (2012). This detection is challenging because of the simultaneous presence of the SGWBs from CBCs, which effectively act as a foreground, limiting the sensitivity to other subdominant SGWBs Callister et al. (2016).
Several methods have been proposed to bypass the CBC foreground. Given the high fraction of individually resolved CBC signals with XG detectors, one could simply fit and subtract them from the data in the frequency domain to lower the foreground Regimbau et al. (2017); Sachdev et al. (2020). However, the recovery of detected signals is never perfect, and the pile-up of residuals from imperfect subtraction can produce an effective foreground that is comparable to the original SGWB before removal Zhou et al. (2023, 2022); Song et al. (2024). Techniques have been proposed to further reduce this effective foreground in the frequency domain, from applying a projection method Cutler and Harms (2006); Sharma and Harms (2020) to estimating and subtracting the expected residual power Pan and Yang (2023). Ref. Belgacem et al. (2024) showed that subtraction results in the frequency domain are actually significantly more optimistic if one abandons the linear-signal approximation in favor of more realistic estimates of the residuals. On the other hand, Refs. Zhong et al. (2023, 2024) proposed a time-frequency approach to notch-out the detected signals by masking the associated pixels. They showed that this method could reduce the CBC foreground by one order of magnitude even when residuals from imperfect recovery are taken into account Zhong et al. (2024).
All of the aforementioned techniques can only lower the CBC foreground by removing individually resolved CBC signals, but the sensitivity to subdominant SGWBs would still be limited by the BNS signals that are too weak to be individually detected, where the unresolvable BNS foreground is about at the order of $\mathcal{O}(10^{-11})$ at $f_{\mathrm{ref}}=25$Hz Zhong et al. (2024); Bellie et al. (2023). Refs. Smith and Thrane (2018); Smith et al. (2020); Biscoveanu et al. (2020) proposed a Bayesian framework to simultaneously fit all of the CBC signals without distinguishing between resolved and unresolved. Ref. Li et al. (2024) suggested using the information on the CBC population from detected events to predict the foreground from unresolved signals and remove it from the data. While promising, the computational feasibility of these methods with a realistic number of events in the context of XG detectors has yet to be proven.
In this Letter, we propose a novel method that combines the notching procedure of Refs. Zhong et al. (2023, 2024) with the joint hierarchical inference of Ref. Callister et al. (2020). First, we notch out all the detected BBH signals, showing that the residual BBH foreground is negligible. Then, we jointly fit within a unified Bayesian framework the resolved BNS events, the foreground from unresolved BNS events, and a CGWB. For illustration, we apply this procedure to mock data including realistic BBH and BNS populations and a flat CGWB.
Astrophysical population and setup. For both BBHs and BNSs, we assume the same population models as in Refs. Zhong et al. (2023, 2024). We employ the PowerLaw+Peak model from the latest LIGO-Virgo-KAGRA (LVK) catalog Abbott et al. (2023) to characterize the BBH mass distribution, while we assume that the BNS masses are uniformly distributed within the range $[1,2]\,\rm{M_{\odot}}$ Abbott et al. (2023); Landry and Read (2021).We assume zero spins and an isotropic distribution of the orbital orientation and sky position for all the binaries. To obtain the redshift distribution, we convolve the star formation rate (SFRFinkel et al. (2022)
$R_{f}(z)=\mathcal{N}\frac{ae^{b(z-z_{p})}}{a-b+be^{a(z-z_{p})}}$ (1)
with a power-law time delay distribution $p(t_{d})\propto t_{d}^{p}$ with $p=-1$. The source-frame merger rate $R_{m}(z)$ then reads
$R_{m}(z)=\int_{t_{\min}}^{t_{\max}}R_{f}(z_{f})p(t_{d})\text{d}t_{d}\,,$ (2)
where $z_{f}=z[t(z)-t_{d}]$ is the binary formation time, we set $t_{\min}=20$ Myr for BNSs and $50$ Myr for BBHs, and we set $t_{\max}$ to be the Hubble time in both cases. We fix the parameters in the SFR to $z_{p}=2.00$, $a=2.37$ and $b=1.80$ Finkel et al. (2022), while $\mathcal{N}$ is a normalization factor chosen so that the local merger rate $\mathcal{R}_{0}$ is consistent with the LVK results Abbott et al. (2023). In particular, for BNSs we set $\mathcal{R}_{0}=320\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$; for BBHs, we choose a value of $\mathcal{R}_{0}$ such that $\mathcal{R}_{m}(z=0.2)=28.1\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ Abbott et al. (2023).
Besides the CBC signals, we also simulate a power-law CGWB
$\Omega_{\mathrm{GW}}(f)=\Omega_{\mathrm{ref}}\left(\frac{f}{f_{\mathrm{ref}}}\right)^{\alpha}\,,$ (3)
where we choose $\alpha=0$ and ${}\Omega_{\mathrm{ref}}=2.0\times 10^{-11}$ at $f_{\mathrm{ref}}=25\,\rm{Hz}$. Our framework is not fundamentally limited by the exact functional form of the CGWB, and we choose a simple flat CGWB for demonstration purposes.
We consider a network of XG interferometers composed of two 40-km long CE detectors located at the LIGO Livingston and LIGO Hanford sites Aasi et al. (2015), and an ET detector located at the Virgo site Acernese et al. (2015). We assume Gaussian and stationary detector noise, modeled via a perfectly known power spectral density (PSD) and uncorrelated among different detectors. For both CE detectors, we adopt the PSD curve of Ref. Gupta et al. (2023), while for ET we assume the PSD of the triangular configuration in Ref. Branchesi et al. (2023a). We set the low-frequency sensitivity limit to be $5~{}\rm{Hz}$ for all detectors in the network.
We use Bilby Ashton et al. (2019) to simulate and add detector noise and CBC signals into a time series. For BBH signals, we adopt the IMRPhenomXAS waveform model Pratten et al. (2020), which is a full inspiral-merger-ringdown, non-precessing waveform model for the fundamental mode. For BNS signals we use the inspiral-only TaylorT4 waveform model Buonanno et al. (2003), as we expect most of the BNS signals in XG detectors to be inspiral-dominated.
We use pygwb Renzini et al. (2023) to simulate and add the CGWB into the time series and compute cross-correlations for the stochastic search. With our network of XG detectors, Ref. Zhong et al. (2024) showed that combining cross-correlation results from all ten possible baselines yields an improvement in the sensitivity of the stochastic search of less than $10\%$ compared to using a single baseline of two CE detectors. Using all baselines would require 10 times more computational power. Therefore, for demonstration purposes, here we only cross-correlate time series data from two CE interferometers.
Data-analysis method. We devise a procedure that combines the notching method of Refs. Zhong et al. (2023, 2024) with the joint Bayesian inference of Ref. Callister et al. (2020).
We assess the detectability of individual CBC signals by computing their network signal-to-noise ratio (SNRSathyaprakash and Schutz (2009), choosing detection thresholds of $\mathrm{SNR}_{\mathrm{thr}}^{\mathrm{BBH}}=8$ for BBHs and $\mathrm{SNR}_{\mathrm{thr}}^{\mathrm{BNS}}=12$ for BNSs Zhong et al. (2024). With the population model we adopt, we find that 99.7% of the BBH and 50.4% of the BNS signals are detected given two CE and one ET detectors network. We then apply the time-frequency approach of Ref. Zhong et al. (2024) to notch out all the detected BBH signals from the data. As we show below, the remaining BBH foreground, resulting from the combination of the few unresolved BBH signals and the residuals from imperfect removal, is weaker than the network sensitivity; hence, we do not include it in the inference studies.
After applying this notching procedure, the data contains only BNS signals and the CGWB. If we followed Ref. Zhong et al. (2024) and notched out the resolved BNS events as well, we would still be limited by the foreground from unresolved BNS signals. Instead, we adopt the Bayesian framework of Ref. Callister et al. (2020) to jointly fit the resolved BNS signals along with the BNS foreground and the CGWB. Ref. Callister et al. (2020) devised this method in the context of current LVK detectors, where the SGWB complements the hierarchical inference on individually resolved events by providing information on the CBC population at high redshift Abbott et al. (2023). Here, instead, the information from detected events helps us tighten the constraints on the BNS foreground and distinguish it from the CGWB.
Let us consider a set of data segments $\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}$ from $N_{\mathrm{obs}}$ individual BNS detections and the cross-correlation spectrum $\hat{C}^{\star}(f)$ after notching out the resolvable BBH signals (see the Supplement for the definition of $\hat{C}^{\star}(f)$). Their joint likelihood can be written as Callister et al. (2020)
($\mathscr{L}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}},\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})=\mathscr{L}_{\mathrm{BNS}}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}|\bm{\Lambda}\bm{}\mathrm{BNS})\times\mathscr{L}_{\mathrm{SGWB}}(\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})\,,$ (4)
where $\mathscr{L}_{\mathrm{BNS}}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}|\bm{\Lambda}\bm{}\mathrm{BNS})$ is the usual hierarchical likelihood for population inference on the detected events Mandel et al. (2019); Taylor and Gerosa (2018), while $\mathscr{L}_{\mathrm{SGWB}}(\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB})$ is the Gaussian likelihood for stochastic searches Mandic et al. (2012); Callister et al. (2017). Here $\bm{\Lambda}\bm{}\mathrm{CGWB}=\{\Omega_{\mathrm{ref}},\alpha\}$ encodes the parameters characterizing the CGWB, while $\bm{\Lambda}\bm{}\mathrm{BNS}$ is the set of hyperparameters defining the BNS population, which affect both the individually detected events and the SGWB from unresolved BNS signals. To simplify the inference, we assume that the BNS mass distribution is perfectly known and that the angular parameters are isotropically distributed. Hence, the only BNS hyperparameters $\bm{\Lambda}\bm{}\mathrm{BNS}$ that we are inferring are the ones characterizing the redshift distribution. In particular, we consider two cases:
  • 1)
    A simplistic known SFR scenario where we assume that the SFR is completely known, and all that is left to determine are the local merger rate, $\mathcal{R}_{0}$, and the power-law index of the time-delay distribution, $p$. Then the joint inference involves a total of four parameters $\{\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB}\}=\{\mathcal{R}_{0},p,\Omega_{\mathrm{ref}},\alpha\}$.
  • 2)
    A more realistic unkown SFR scenario where we assume no prior knowledge of the redshift distribution, except for its functional form. Following the approach of Ref. Ng et al. (2021), we assume that the merger rate has the same functional form as the SFR of Eq. (1), then we fit this function to the merger rate $R_{m}(z)$ computed via Eq. (2) to obtain the injected values $a_{\mathrm{inj}}=2.15,~{}b_{\mathrm{inj}}=1.40,~{}z_{p}^{\mathrm{inj}}=1.64$. In this case, we directly infer these parameters along with the local merger rate $\mathcal{R}_{0}$, so the joint inference involves a total of six parameters $\{\bm{\Lambda}\bm{}\mathrm{BNS},\bm{\Lambda}\bm{}\mathrm{CGWB}\}=\{\mathcal{R}_{0},a,b,z_{p},\Omega_{\mathrm{ref}},\alpha\}$.
Performing full Bayesian parameter estimation (PE) on thousands of individually resolved BNS signals is currently computationally unfeasible. Hence, we generate synthetic samples for the individual BNS likelihoods on source parameters using the approximate method of Refs. Fishbach et al. (2018); Callister et al. (2020). For the exact expressions of the joint likelihood and additional details on the inference procedure, we refer the reader to the Supplemental Material.
Joint inference results. For demonstration purposes, we perform our analysis assuming an observation time $T_{\mathrm{obs}}=10$ days. With our population model, this corresponds to roughly $\sim 650$ BBH signals that we notch out, and $\sim 7000$ resolvable BNS events that enter the hierarchical likelihood $\mathscr{L}_{\mathrm{BNS}}(\{\bm{d}_{i}\}_{i=1}^{N_{\mathrm{obs}}}|\bm{\Lambda}\bm{}\mathrm{BNS})$.
In vanilla searches for a single, dominant SGWB, its SNR is straightforward to compute, and it is a good figure of merit to estimate whether the target SGWB is strong enough to be detectable relative to the detector noise Abbott et al. (2023); Allen and Romano (1999); Abbott et al. (2021). In our case, instead, we are jointly fitting for the BNS hyperparameters from individually detected sources plus two distinct SGWBs, and our target is, in fact, the subdominant CGWB. Hence, we turn to a full Bayesian approach and we study the inferred posterior to find evidence for CGWB detection.
Figure 1: Marginalized Joint posterior distributions on $\{\mathcal{R}_{0},\alpha,\Omega_{\mathrm{ref}}\}$ for the models assuming known SFR (purple curves) and unknown SFR (green curves). Contours enclose $68\%$ and $95\%$ of the probability mass. Gray dashed lines indicate the true simulated values of parameters.
In both our scenarios, we find that the posteriors on all the parameters are fully consistent with the simulated values. In Fig. 1, we compare the joint posteriors on the common parameters for both inferences, namely the CGWB parameters $\{\Omega_{\mathrm{ref}},\alpha\}$ and the local BNS merger rate $\mathcal{R}_{0}$. As expected, we find that assuming better knowledge of the BNS redshift distribution has a large impact on inferring the local merger rate. The $95\%$ confidence interval on $\mathcal{R}_{0}$ roughly doubles going from known (purple curves) to unknown (green curves) SFR.
However, the joint posterior on $\{\alpha,~{}\Omega_{\mathrm{ref}}\}$ is remarkably similar in both analyses, indicating that adding more parameters to our inference of the BNS redshift distribution does not have a large impact on the inference of the CGWB. Crucially, we find that we can confidently claim the detection of the CGWB in both scenarios: we exclude $\Omega_{\mathrm{ref}}=0$ at the $6.8\,\sigma$ level assuming known SFR and at $6.1\,\sigma$ level with unknown SFR. We refer to Table 1 for the marginalized posterior distributions of all parameters in the different cases.
Table 1: Marginalized posterior distribution (median and 95% credible symmetric interval).
Scenario $\mathcal{R}_{0}$ $p$ $a$ $b$ $z_{p}$ $\Omega_{\mathrm{ref}}$ $\alpha$
Joint search: Known SFR $319^{+25}_{-24}$ $-1.00^{+0.04}_{-0.04}$ $1.92^{+0.58}_{-0.52}\times 10^{-11}$ $-0.10^{+0.32}_{-0.37}$
Joint search: Unknown SFR $305^{+48}_{-47}$ $2.24^{+0.19}_{-0.16}$ $1.43^{+0.28}_{-0.24}$ $1.64^{+0.07}_{-0.07}$ $2.04^{+0.66}_{-0.66}\times 10^{-11}$ $-0.06^{+0.31}_{-0.39}$
Ideal search - - - - - $1.95^{+0.35}_{-0.34}\times 10^{-11}$ $-0.14^{+0.29}_{-0.31}$
Figure 2: Joint posterior distributions on $\{\Omega_{\mathrm{ref}},\alpha\}$ across three different scenarios: known SFR (purple curves), unknown SFR (orange curves), and an ideal search without any CBC foreground (green curves). Contours represent $68\%$ and $95\%$ of the probability.
In Fig. 2, we further compare the marginalized posterior distributions on $\{\Omega_{\mathrm{ref}},\alpha\}$ from our two joint-inference scenarios (purple and green curves) with the posterior from an ideal search for the CGWB where no CBC foreground is present (orange curves). In other words, the ideal search corresponds to a benchmark case where only the CGWB is simulated, and the second term of Eq. (4) in the form $\mathscr{L}_{\mathrm{SGWB}}(\hat{C}^{\star}(f)|\bm{\Lambda}\bm{}\mathrm{CGWB})$ is used for the inference. After applying our notching+joint inference scheme, the recovery of the CGWB is not significantly affected by the presence of a CBC foreground. We find that the $1\,\sigma$ error on the CGWB amplitude $\Omega_{\mathrm{ref}}$ increases by a factor of $\sim 90\%$ ($\sim 60\%$) in the unknown (known) SFR case compared to the ideal search with no foreground present. The difference in the marginalized posterior for the power-law index $\alpha$ is even milder, with a 19% (17%) increase in the $1\,\sigma$ error for the unknown (known) SFR scenario with respect to the ideal search.
Figure 3: The sensitivity of the CGWB search. The top four dashed curves show the baseline sensitivity to stochastic searches given different observation times: ten days (pink curve), one month (blue curve), one year (green curve), and ten years (orange curve). The straight bands represent the uncertainty on the BNS foreground for each observation time, with the same color scheme. Within each band, the upper (lower) line corresponds to the unknown (known) SFR scenario. The gray curve is the remaining BBH foreground after notching with 10 days of observation.
Sensitivity of the CGWB search. In order to further assess the constraining power of our procedure, in Fig. 3 we compare the noise level for stochastic searches, the uncertainty on the BNS foreground, and the residual BBH foreground after notching out the resolved signals. The baseline noise level $\sigma_{n}(f)$ is computed through Eqs. (S1) and (S2) of the Supplement. The uncertainty on the inferred BNS foreground is estimated by computing energy spectra $\Omega_{\mathrm{BNS}}(f)$ for all different posterior samples and the resulting standard deviation $\sigma_{\Omega_{\mathrm{BNS}}}(f)$ at each frequency (see the Supplement for the description of $\Omega_{\mathrm{BNS}}(f)$ calculation). The sensitivity of stochastic searches scales as $\sim 1/\sqrt{T_{\mathrm{obs}}}$ Allen and Romano (1999), so we show the comparison for different values of $T_{\mathrm{obs}}$ by simply rescaling $\sigma_{\Omega_{\mathrm{BNS}}}^{\mathrm{T^{\prime}_{obs}}}(f)=\sigma_{\Omega_{\mathrm{BNS}}}^{\mathrm{T_{obs}}}(f)\left(\sqrt{T_{\mathrm{obs}}}/\sqrt{T^{\prime}_{\mathrm{obs}}}\right)$. We also explicitly check this scaling by running the same full simulation using fewer data, in which case we do observe the width of the posterior distributions of parameters scale as $1/\sqrt{T_{\mathrm{obs}}}$ that further makes $\sigma_{\Omega_{\mathrm{BNS}}}\sim 1/\sqrt{T_{\mathrm{obs}}}$
By only injecting BBHs signals into the time series and notching out the resolvable ones in the time-frequency domain as we did in Ref. Zhong et al. (2023, 2024), we find that the residual BBH foreground is $\sim 100$ times lower than the baseline noise $\sigma_{n}(f)$ at every frequency, even for 10 yrs observation, which justifies neglecting its contribution to the joint likelihood in Eq. (4). The uncertainty on the BNS foreground differs by about 17% between the inference with known and unknown SFR, and it is well below the baseline noise at all frequencies, for both known and unknown SFR, and for all observation times. This is consistent with the results discussed in the previous section: the BNS foreground is well constrained in both our inferences, allowing us to decouple it from the CGWB. By applying the $\sim 1/\sqrt{T_{\mathrm{obs}}}$ scaling on the CGWB uncertainty, we can project sensitivity estimates for longer observations than 10 days. For the flat CGWB, we find that we can exclude $\Omega_{\mathrm{ref}}=0$ at $5\,\sigma$ level for $\Omega_{\mathrm{ref}}=2.7\times 10^{-12}/\sqrt{T_{\mathrm{obs}}}$ ($2.5\times 10^{-12}/\sqrt{T_{\mathrm{obs}}}$) in the unknown (known) SFR case, where $T_{\mathrm{obs}}$ is the observation time in years.
Conclusions. In this work, we combine the notching procedure of Refs. Zhong et al. (2023, 2024) with the joint-likelihood approach of Ref. Callister et al. (2020) to search for CGWBs with XG detectors. In this context, we demonstrate a methodology capable, for the first time, of simultaneously handling BBH signals, BNS signals, and a CGWB. Considering a simple population model for the hierarchical inference on the detected BNS events and a flat CGWB, we find that we can successfully claim detection of a sufficiently loud CGWB within 10 days of observation. In particular, for a flat CGWB with amplitude of $\Omega_{\mathrm{ref}}=2.0\times 10^{-11}$, we find that we can exclude $\Omega_{\mathrm{ref}}=0$ at at least $\sim 6\,\sigma$ confidence level, depending on the assumptions on the BNS population model. The uncertainty on the recovered BNS foreground is well below the detector noise level. If we project the sensitivity to 1 year of observation, then we will reach the $5\sigma$ sensitivity of $\Omega_{\mathrm{ref}}=2.7\times 10^{-12}$, which is within a factor of $\lesssim 2$ from the ideal stochastic search in absence of CBC foregrounds.
This study represents a crucial step forward in developing methods to detect CGWBs with XG detectors, but more work is needed before such methods can be applied to real data:
(1) We have assumed that the detector noise is perfectly known, while in practice, it can be challenging to estimate the PSD in the presence of glitches, non-stationary noise, cosmological and astrophysical SGWBs.
(2) We have neglected the contribution of NSBHs, which can give rise to an unresolved foreground that is comparable to the one from BNSs in XG detectors Bellie et al. (2023). Our formalism can be straightforwardly extended to include a NSBH population. However, this would inevitably increase the computational cost and it could worsen the constraints on some of the inferred parameters.
(3) When performing hierarchical inference on the resolved BNS signals, we generate mock likelihood samples on the source parameters of individual events by using a Newtonian approximation of the signals and a Gaussian approximation of the likelihood Fishbach et al. (2018); Callister et al. (2020). This method can underestimate errors on source parameters compared to performing full Bayesian PE with more realistic waveform models, although we still expect the population hyperparameters to be tightly constrained when the number of events is large enough De Renzis et al. (2024). Sampling individual-event likelihoods with full PE and state-of-the-art waveform models for several thousands of BNS signals in XG detectors is currently computationally prohibitive. Several promising techniques to speed up PE in preparation for XG detectors are being developed Green and Gair (2021); Dax et al. (2023); Wong et al. (2023); Alvey et al. (2023) and can be incorporated in our formalism in the future.
(4) We assume that the detected CBC signals originate from the same astrophysical population that generates the CBC foreground from unresolved signals. The presence of an unmodeled sub-population predominantly at high redshifts (e.g., from Population III stars Martinovic et al. (2022) or primordial black holes Mukherjee et al. (2022)) could potentially give rise to biases in the estimate of the CBC foreground, and thus affect the detection of the CGWB.
We leave the exploration of these effects to future work.
Acknowledgements We thank Will M. Farr, Thomas Callister, and Francesco Iacovelli for helpful discussions. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation (NSF) Grants PHY-0757058 and PHY-0823459. B.Z. is supported by the Fermi Research Alliance, LLC, acting under Contract No. DE-AC02-07CH11359. E.B. and L.R. are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation grant No. PGR01167. H.Z. and V.M. are in part supported by the NSF grant PHY-2409173. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103.